# Sparsing Law: Towards Large Language Models with Greater Activation Sparsity

### v1

@@ 5 Discussion
In this section, we mainly discuss how this work can help accomplish our key target: how to obtain an LLM with greater activation sparsity.
Architectural design
Before kicking off the pre-training stage, a wiser design of the LLM architecture can effectively improve the limit activation sparsity. Based on the discoveries in this paper, we propose to replace SiLU with ReLU as the activation function to promote greater sparsity and leverage its increasing trend of sparsity with more training data. Moreover, a deeper model can potentially exhibit higher sparsity given a fixed computation budget. Nevertheless, as a model with an extreme width-depth ratio can violate the performance, the best setting for the width and depth still deserves careful studies, combined with the quantitative study on the relationship between performance and the width-depth ratio.
Training-time predictable sparsity
Another important value of our work lies in the prediction of sparsity during the pre-training stage. By fitting the power-law or logspace power-law between activation sparsity and the amount of training data, model developers can either predict the theoretical upper/lower bound sparsity of a model to evaluate its potential (e.g., in inference acceleration), or estimate the number of tokens required to achieve a desired sparsity ratio.
Lens for the convergence of neuron specialization
Generally, the loss curve is one of the most important signs of the training state (e.g., at which point the model converges). However, if we compare the loss curve in Figure 12 and the sparsity (activation ratio) curve in Figure 4, we will find that the convergence speed of activation sparsity is much slower than loss, indicating the ongoing of neuron specialization even when the training loss changes little. Despite the wide recognition of the neuron specialization phenomenon (Li et al., 2022; Zhang et al., 2023), it is still unclear when such specialization converges and how to inspect this progress. Besides, the loss curve is often not a good standard for convergence, especially for modern LLMs with trillion-level pre-training data. We believe that the trend of activation sparsity can provide a lens for inspecting the progress of neuron specialization as well as the training convergence. We leave the study on the correlation between activation sparsity and neuron specialization for future work.
@@ 6 Conclusion
In this work, we first propose a precise, versatile, and performance-aware metric for activation sparsity, called PPL-$p\%$ sparsity, whose rationality is demonstrated in striking a better trade-off between performance and sparsity. Next, we conduct a comprehensive and quantitative study on how the activation sparsity scales with the amount of training data and parameter scales. The influence of two key architectural settings, namely the activation function and the width-depth ratio, are also well evaluated. Through extensive experiments, we figure out the quantitative sparsity-data relationship, substantiate the advantage of ReLU activation, find a small width-depth ratio helpful in promoting sparsity, observe and then explain the insensitiveness of sparsity to scale. These can better instruct LLM developers to build models with greater activation sparsity and leverage the merits of sparsity.
Limitations
One limitation lies in the lack of experiments on even larger LLMs, such as models with more than 7B parameters. Such experiments will cost considerable computation resources and time but may potentially encounter unexpected observations (e.g., emergent properties).
@@
Another drawback of our study is the absence of computation (e.g., FLOPS) in some analyses, especially the experiments for width-depth ratios. In Section 4.3, we find a smaller width-depth ratio potentially produces a sparser model. However, with the substantial increase in the number of layers, the training efficiency is significantly decreased, as we have observed in the training process. Therefore, in addition to performance, the computation costs of a model also deserve consideration. Similarly, considering the values of activation sparsity in acceleration, it may be interesting to involve the inference computation as a variable in our study.
@@
Finally, an obvious limitation of our PPL-$p\%$ metric (as well as all the sparsity metrics relying on a validation dataset) is the sensitivity to different data distributions. Intuitively, the same model can have different sparsity ratios on distinct datasets or tasks. The correlation between sparsity and influential factors (e.g., the form of power laws) can also have dataset-unique characteristics. A piece of evidence already presented in our paper is in Table 1, where the performance on commonsense reasoning tasks is insensitive to $p\%$, largely different from the results on reading comprehension tasks. Moreover, the data mixing policies for pre-training can also have a considerable impact on the activation sparsity, which we leave for future work.
Appendix A Training Loss Dynamics
@@
To present the comprehensive training dynamics of our pre-trained models, we plot the trend of loss with the increase of training data in Figure 12. As can be clearly observed, larger models have smaller training loss. Besides, we also plot the limit values of the training loss with infinite training tokens, shown in Figure 12. As demonstrated in the above two figures, SiLU and ReLU models are well comparable from the loss aspect.
@@
Figure 12: The limits of the training loss with the amount of training data approaches infinity.
Appendix B Sparsity Stabilizing Strategy
@@
We find that the stochastic factors in gradient descent during pre-training have a significant impact on the metric of activation sparsity. Especially, during the early training stage, the model is far from convergence with considerable sparsity fluctuations, and the magnitude-based sparsity metric can become unstable, making the sparsity-data curve not smooth enough.
@@
To eliminate the influence of these unstable factors to facilitate smoother sparsity metric, we first drop the sparsity points during the warmup stage for curve fitting. Moreover, we mainly apply the PPL-$p\%$ on the last several checkpoints (specifically, the last five pre-trained checkpoints) as a whole, binary-searching a CETT value that controls the average PPL of these checkpoints to just increase by $p\%$. Then this CETT value is applied to all the checkpoints of this pre-training process to measure the sparsity.
Appendix C Binary Search Algorithm for CETT
@@
Given a list of checkpoints, a validation dataset, and a hyper-parameter $p\%$, we employ Algorithm 1 to find the CETT value that just makes the average PPL of these checkpoints on the validation dataset rise by exactly $p\%$, compared to the dense setting with all the neurons activated. Note that this algorithm can be applied to either a single checkpoint or multiple checkpoints, as adopted in the strategy described in Appendix B.
Algorithm 1 Find the CETT value for PPL-$p\%$ sparsity
@@
  Input: The hyper-parameter $p\%$.
  Input: The error tolerance $eps$.
  Output: The $CETT$ that just makes the average PPL of $CkptList$ on $ValidSet$ rise by $p\%$.
@@
Appendix D Fitting Algorithm and Results
We employ the Levenberg-Marquardt method (Marquardt, 1963) to fit the activation-data curves. To improve the stability of curve fitting, we divide the number of tokens passed (i.e., the amount of training data) by $10^{9}$ to normalize its magnitude. All the results we obtained from fitting Eq. (4) (for ReLU-activated models) and Eq. (5) (for SiLU-activated models) are shown in Table 2.
$\alpha$ $b$ $c$ $A_{0}$
SiLU 0.1B $4.79\times 10^{-01}$ $1.02\times 10^{-01}$ $4.09\times 10^{-01}$
0.2B $8.44\times 10^{-01}$ $2.08\times 10^{-01}$ $3.90\times 10^{-01}$
0.4B $1.03\times 10^{+00}$ $4.20\times 10^{-01}$ $3.85\times 10^{-01}$
0.8B $9.95\times 10^{-01}$ $5.62\times 10^{-01}$ $3.83\times 10^{-01}$
1.2B $9.67\times 10^{-01}$ $5.38\times 10^{-01}$ $3.82\times 10^{-01}$
Appendix E Datasets and Benchmarks
Training data
Validation data
To measure the PPL-$p\%$ sparsity more precisely, we introduce a tiny validation dataset, which shares the same distribution as the pre-training data. We conduct deduplication to eliminate any intersections between validation and pre-training data.
Evaluation benchmarks
Appendix F Detailed Training Settings
We utilize the MiniCPM (Hu et al., 2024) architecture and adopt its hyper-parameter policies, along with the WSD learning rate scheduling method. Across all parameter scales, the ratio of $d_{f}$ to $d_{h}$ equal to 2.5 consistently, the number of query heads always matches that of key and value heads, and the width-depth ratios range from 48 to 56, generally similar across different scales. The specific number of parameters of various settings are shown in Table 3. We employ the following pre-training hyper-parameters across all settings: peak learning rate $lr=0.01$, $\beta_{1}=0.9$, $\beta_{2}=0.95$, $weight\ decay=0.1$. The batch size depends on the parameter scale, as presented in Table 3.
Table 3: Hyper-parameters across various parameter scales.
Parameter Scale 0.1B 0.2B 0.4B 0.8B 1.2B
##\## non-embedding parameters $1.08\times 10^{8}$ $2.41\times 10^{8}$ $4.52\times 10^{8}$ $7.60\times 10^{8}$ $1.18\times 10^{9}$
batch size $3.27\times 10^{5}$ $5.90\times 10^{5}$ $7.86\times 10^{5}$ $1.18\times 10^{6}$ $1.57\times 10^{6}$
Appendix G Dataset-wise Activation Pattern
Although the overall distribution patterns of activation frequencies are similar in terms of the average scenario, they exhibit difference when focusing on neurons in specific layers, such as the first, the last, or the exact middle layer. As shown in Figure 13, models with varying parameter scales have diverse neuron activation frequency distributions in the first layer and the last layer, while the patterns on the middle layer are still largely scale-insensitive.
@@ Appendix G Dataset-wise Activation Pattern
Appendix H Performance on Independent Benchmarks
In Table 1, we already provide the average performance on the two groups of commonsense reasoning and reading comprehension. In this section, we present the evaluation scores on independent benchmarks of these two task groups, as shown in Table 4 and Table 5, respectively. From these tables, it can be observed that in commonsense reasoning benchmarks, as the number of model parameters increases from 0.1B to 1.2B, the average evaluation score of the ReLU settings rises from 49.6 to 60.0, while the average score of the SiLU settings increases from 49.5 to 59.6. Similarly, in reading comprehension benchmarks, the score of ReLU settings goes from 28.2 to 53.2, and the score of SiLU settings rises from 27.7 to 54.8. Additionally, models with these two distinct activation functions demonstrate comparable performance at the same parameter scale. Moreover, under the PPL-1% setting, the models are generally comparable to the dense setting with all neurons activated, whereas under the PPL-5% setting, they tend to suffer from significant performance on reading comprehension tasks, but the commonsense reasoning scores almost remain unaffected, which is a phenomenon worth studies.

### v2

@@ 7 Discussion
In this section, we mainly discuss other potential values of this work, especially in terms of monitoring the model dynamics from new aspects.
Training-time predictable sparsity Our work enables the prediction of sparsity during the training stage. By fitting the power-law or logspace power-law between activation sparsity and the amount of pre-training data, model developers can either predict the theoretical upper/lower bound sparsity of a model to evaluate its potential (e.g., in inference acceleration), or estimate the number of tokens required to achieve a desired sparsity ratio.
Lens for the convergence of neuron specialization Generally, the loss curve is an important signal of the training state (e.g., at which point the model converges). However, if we compare the loss curve in Figure 14 and the sparsity (activation ratio) curve in Figure 4, we will find that the convergence speed of activation sparsity is much slower than loss, indicating the ongoing of neuron specialization even when the loss changes little. Despite the wide recognition of the neuron specialization phenomenon (Li et al., 2022; Zhang et al., 2023), it is unclear when such specialization converges and how to inspect this progress. Besides, the loss curve is often not a good inspector for convergence, especially for LLMs with trillion-level pre-training data. We believe that the trend of activation sparsity can provide a new lens for inspecting the progress of neuron specialization and training convergence.
@@ 8 Conclusion
In this work, we address three important research questions about activation sparsity. First, we prove CETT-PPL-1% as a good sparsity metric with the most accurate recognition of weakly-contributed neurons and negligible performance degradation. Next, quantitative studies are conducted on the influential factors of activation sparsity. Finally, towards a sparser LLM, our findings substantiate the advantage of ReLU activation, more training data, and a small width-depth ratio. We also observe and explain the scale insensitivity of sparsity. These can better instruct LLM developers to build sparser LLMs and leverage their merits.
Impact Statement
This paper presents work whose goal is to advance the field of large language models and address underexplored issues of activation sparsity. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here. For more rigorous studies, we also discuss the limitations of our work in Appendix A.
Appendix A Limitations
@@
One drawback of our study is the absence of computation (e.g., FLOPS) in some analyses, especially the experiments for width-depth ratios. In Section 5.2, we find a smaller width-depth ratio potentially produces a sparser model. However, with the substantial increase in the number of layers, the training efficiency is significantly decreased, as we have observed in the training process. Therefore, in addition to performance, the computation costs of a model also deserve consideration. Similarly, considering the values of activation sparsity in acceleration, it may be interesting to involve the inference computation as a variable in our study.
@@
Another limitation of our CETT-PPL-$p\%$ metric (as well as all the sparsity metrics relying on a validation dataset) is the sensitivity to different data distributions. Intuitively, the same model can have different sparsity ratios on distinct datasets or tasks. The correlation between sparsity and influential factors (e.g., the form of power laws) can also have dataset-unique characteristics. A piece of evidence already presented in our paper is in Table 1, where the performance on commonsense reasoning tasks is insensitive to $p\%$, largely different from the results on reading comprehension tasks. Moreover, the data mixing policies for pre-training can also have a considerable impact on the activation sparsity, which we leave for future work.
Appendix B Activation Sparsity and Mixture-of-Experts
Mixture-of-experts (MoE) is a mainstream method to achieve high activation sparsity by introducing constraints at the model architecture level. Typically, MoE uses a token-level top-k parameter selection router to assign a fixed sparsity ratio for each token at each layer (Fedus et al., 2022; Zoph et al., 2022). However, these constraints often sacrifice model flexibility and performance. Recent works reveal the potential performance degradation caused by such inflexible sparsity assignment (Huang et al., 2024; Liu et al., 2024). Moreover, to inspect the impact of such constraints, we plot the PPL-activation (PPL denotes perplexity) Pareto curve of MoE in Figure 12 and compare it with a vanilla decoder-only Transformer (Touvron et al., 2023) of the same parameter scale and amount of pre-training data [^3] MoE models of different sparsity are obtained by tuning the number of activated experts, while for the vanilla setting, we adjust the CETT hyper-parameter proposed by Zhang et al. (2024a).. MoE has a significantly worse performance-sparsity trade-off. Moreover, the best sparsity ratio is also hard to predefine, since a too-high or too-low sparsity ratio may lead to more severe performance degradation or substantial unnecessary computation, respectively.
Therefore, to avoid negative impacts on flexibility and performance, in this work, we focus on the intrinsic activation sparsity within vanilla decoder-only Transformer-based LLMs in this paper, such as GPT (Brown et al., 2020) and LLaMA (Touvron et al., 2023).
Appendix C Activation Sparsity and Weight Pruning
Weight pruning, also a popular paradigm for inference acceleration, accelerates LLMs by removing certain elements from the model parameters (e.g., neurons, weights, or structured blocks). Nevertheless, the sparsity introduced by weight pruning is fundamentally different from activation sparsity, as the pruning sparsity is completely static, namely independent of inputs. Specifically, weight pruning always drops the same part of parameters regardless of inputs. Consequently, enforcing high such static sparsity can easily cause considerable performance degradation (Frantar & Alistarh, 2023; Xia et al., 2023).
By contrast, activation sparsity dynamically determines weakly-contributed neurons given specific input tokens, which potentially sacrifices less LLM capacity and performance. For example, ReLU-activated models, with considerably higher activation sparsity, have comparable performance to mainstream SiLU-activated ones. Moreover, weight pruning and activation sparsity can also be combined to further promote LLM efficiency.
Appendix D Training Loss Dynamics
@@
To present the comprehensive training dynamics, we plot the trend of loss with the increase of training data in Figure 14. As can be clearly observed, larger models have smaller training loss. Besides, we also plot the limit values of the training loss with infinite training tokens, shown in Figure 14. As demonstrated in the above two figures, SiLU and ReLU models are well comparable from the loss aspect.
@@
Figure 14: The limits of the training loss with the amount of pre-training data approaches infinity.
Appendix E Sparsity Stabilizing Strategy
@@
We find that the stochastic factors in gradient descent during pre-training have a significant impact on the metric of activation sparsity. Especially, during the early training stage, the model is far from convergence with considerable sparsity fluctuations, and the magnitude-based sparsity metric can become unstable and cause problems to curve fitting.
@@
To eliminate the influence of these unstable factors to facilitate a smoother sparsity metric, we first drop the sparsity points during the warmup stage for curve fitting. Moreover, we mainly apply the CETT-PPL-$p\%$ on the last several checkpoints (specifically, the last five pre-trained checkpoints) as a whole, binary-searching the hyper-parameter value that controls the average PPL of these checkpoints to just increase by $p\%$. Then this hyper-parameter is applied to all the checkpoints of this pre-training process to measure the sparsity ratio.
Appendix F Binary Search Algorithm for CETT
Figure 15: Experiments show that both PPL and the activation ratio change monotonously with CETT, the key hyper-parameter of CETT-PPL-$p\%$. This provides rationality for using a binary search to determine the CETT hyper-parameter.
@@ Appendix F Binary Search Algorithm for CETT
Given a list of checkpoints, a validation dataset, and a hyper-parameter $p\%$, we employ Algorithm 1 to find the CETT hyper-parameter that just makes the average PPL of these checkpoints on the validation dataset rise by exactly $p\%$, compared to the dense setting with all the neurons activated. The rationality of using binary search is substantiated by the monotonous relationship between PPL and CETT, as shown in Figure 15. Note that this algorithm can be applied to either a single checkpoint or multiple checkpoints, as adopted in the strategy described in Appendix E.
Appendix G Fitting Algorithm and Results
We employ the Levenberg-Marquardt method (Marquardt, 1963) to fit the activation-data curves. To improve the stability of curve fitting, we divide the number of tokens passed (i.e., the amount of pre-training data) by $10^{9}$ to normalize its magnitude. All the results we obtained from fitting Eq. (4) (for ReLU-activated models) and Eq. (5) (for SiLU-activated models) are shown in Table 2.
$\alpha$ $b$ $c$ $A_{0}$
2.4B $1.53\times 10^{+00}$ $-3.46\times 10^{+00}$ $1.56\times 10^{-04}$ $6.48\times 10^{-02}$
SiLU 0.1B $4.79\times 10^{-01}$ $1.02\times 10^{-01}$ $4.09\times 10^{-01}$
0.2B $8.44\times 10^{-01}$ $2.08\times 10^{-01}$ $3.90\times 10^{-01}$
0.4B $1.03\times 10^{+00}$ $4.20\times 10^{-01}$ $3.85\times 10^{-01}$
0.8B $9.95\times 10^{-01}$ $5.62\times 10^{-01}$ $3.83\times 10^{-01}$
1.2B $9.67\times 10^{-01}$ $5.38\times 10^{-01}$ $3.82\times 10^{-01}$
Table 3: Hyper-parameters across various parameter scales.
Parameter Scale 0.1B 0.2B 0.4B 0.8B 1.2B 2.4B
##\## non-embedding parameters $1.08\times 10^{8}$ $2.41\times 10^{8}$ $4.52\times 10^{8}$ $7.60\times 10^{8}$ $1.18\times 10^{9}$ $2.44\times 10^{9}$
batch size $3.27\times 10^{5}$ $5.90\times 10^{5}$ $7.86\times 10^{5}$ $1.18\times 10^{6}$ $1.57\times 10^{6}$ $2.10\times 10^{6}$
Appendix H Datasets and Benchmarks
Training data
Validation data
To measure the CETT-PPL-$p\%$ sparsity more precisely, we introduce a tiny validation dataset, which shares the same distribution as the pre-training data. We conduct deduplication to eliminate any intersections between validation and pre-training data.
Evaluation benchmarks
Appendix I Detailed Training Settings
We utilize the $\mu P$ Transformer (Hu et al., 2024) architecture and adopt its hyper-parameter policies, along with the WSD learning rate scheduling method. Across all parameter scales, the ratio of $d_{f}$ to $d_{h}$ is equal to 2.5 consistently, the number of query heads always matches that of key and value heads, and the width-depth ratios range from 48 to 56, generally similar across different scales. The specific number of parameters of various settings are shown in Table 3. We employ the following pre-training hyper-parameters across all settings: peak learning rate $lr=0.01$, $\beta_{1}=0.9$, $\beta_{2}=0.95$, $weight\ decay=0.1$. The batch size depends on the parameter scale, as presented in Table 3.
Appendix J Dataset-wise Activation Pattern
Although the overall distribution patterns of activation frequencies are similar in terms of the average scenario, they exhibit differences when focusing on neurons in specific layers, such as the first, the last, or the exact middle layer. As shown in Figure 16, models with varying parameter scales have diverse neuron activation frequency distributions in the first layer and the last layer, while the patterns on the middle layer are still largely scale-insensitive.
Appendix K Performance on Independent Benchmarks
In Table 1, we already provide the average performance on the two groups of commonsense reasoning and reading comprehension. In this section, we present the evaluation scores on independent benchmarks of these two task groups, as shown in Table 4 and Table 5, respectively. From these tables, it can be observed that in commonsense reasoning benchmarks, as the number of model parameters increases from 0.1B to 1.2B, the average evaluation score of the ReLU settings rises from 49.6 to 60.0, while the average score of the SiLU settings increases from 49.5 to 59.6. Similarly, in reading comprehension benchmarks, the score of ReLU settings goes from 28.2 to 53.2, and the score of SiLU settings rises from 27.7 to 54.8. Additionally, models with these two distinct activation functions demonstrate comparable performance at the same parameter scale. Moreover, under the CETT-PPL-1% setting, the models are generally comparable to the dense setting with all neurons activated, whereas under the CETT-PPL-5% setting, they tend to suffer from significant performance on reading comprehension tasks, but the commonsense reasoning scores almost remain unaffected, which is a phenomenon worth studies.
Appendix L Details of Acceleration Experiments
To demonstrate the practical inference acceleration values of activation sparsity, we run experiments with the 2.4B ReLU-activated model on two different acceleration frameworks: PowerInfer (Song et al., 2023) and “llama.cpp” (Gerganov, 2023). Specifically, PowerInfer, tailored for activation sparsity, involves an offline profiler and online activation predictors to forecast the activation pattern of each neuron. Therefore, PowerInfer can wisely allocate hardware resources according to the activation frequencies of different neurons and save redundant computation and time wasted on weakly-contributed neurons. By contrast, “llama.cpp” does not utilize activation sparsity for acceleration, computing the FFNs in a dense manner.
Both frameworks are compiled with CUDA enabled and run on the same machine with 104 CPUs and 1 NVIDIA A800 GPU. Although “llama.cpp” does not support ReLU and thus cannot correctly conduct inference with our 2.4B model, this does not impact the acceleration experiment as the FLOPS remain the same as a SiLU-activated model. We use 100 test prompts sampled from C4444https://huggingface.co/datasets/allenai/c4, and each prompt is composed of 5 prefix tokens.
Consequently, we found that PowerInfer can perform decoding at an average speed of 41.79 tokens per second, while “llama.cpp” can only reach 10.23 tokens per second. The 4.1$\times$ speedup of PowerInfer provides strong evidence of the acceleration potential offered by activation sparsity.
Algorithm 1 Find the CETT hyper-parameter for CETT-PPL-$p\%$ sparsity
@@ Appendix I Detailed Training Settings
  Input: The PPL increase tolerance $p\%$.
  Input: The error tolerance $eps$ for binary search.
  Output: The hyper-parameter $CETT$ that just makes the average PPL of $CkptList$ on $ValidSet$ rise by $p\%$.